OpenStudy (anonymous):

Which best describes the graph of y = 76(1.013)^x ? a. circle b. parabola c. exponential growth d. exponential decay

OpenStudy (anonymous):

d

OpenStudy (dumbcow):

its an exponential function with the base being greater than 1 ...so exponential growth

OpenStudy (anonymous):

So C or D?

OpenStudy (dumbcow):

a decay function has base less than 1 Ex: y = 0.4^x a circle has the form x^2 + y^2 = r^2 parabola has the form y = ax^2 +bx + c

OpenStudy (anonymous):

So it's C right?

OpenStudy (dumbcow):

yep